# Wealth Thermalization Hypothesis

### v2

@@ Abstract
First, we present in Fig. S9 the Lorenz curve for the data of 504 S&P500 companies of the New York Stock Exchange (NYSE) of June 16, 2025 (see Ref. [33]). This Fig. S9 shows the direct comparison of the Lorenz curve of NYSE and the corresponding RJ thermal distribution of the RJS model (at same Gini value). Here, we use the standard value $N=10000$ for the RJS curve but using a reduced value $N=504$ (as the number of companies) gives the same RJS curve within graphical precision. The quality of agreement with the RJS model is comparable to the cases of US or World in Fig. 3. We also note the characteristic values: $h=0.191$ at wealth $w=0.02$; $w=0.092$ at $h=0.5$; the wealth of top 10 percent of $h$ companies is $0.6021-w(0.9)=$ 0.602 and the wealth of top 1 percent of companies is $0.2671-w(0.99)=$ 0.267. Thus we see that there is a small fraction of oligarchic companies that monopolize a big fraction of total wealth. The fraction of poor companies, corresponding to the RJ condensate, is smaller than the fraction of poor households in the US or World cases. We attribute this to the fact that these 504 companies of S&P500 represent only about 80 percent of the total capitalization of US companies.
@@
Figure S9: Comparison of the Lorenz curve for the S&P500 companies of NYSE 2025 (black; data from Ref. [33]) with the corresponding curve for the RJS model (red curve; $N=10000$) at same Gini coefficient $G=0.692$ obtained for $\varepsilon=0.1582$.
@@
Fig. S10, compares the Lorenz curve for the London stock exchange (2024; data from Ref. [34]) with the RJS model. Here, the Gini coefficient $G=0.9126$ is higher than for the US and World cases and the corresponding value $\varepsilon\approx 0.044$ for the RJS model is quite small. Due to the high value of $G$ the first probability $\rho_{0}=0.6545$ is very high indicating a strong RJ condensation with exactly $w=0$ for $h\in[0,\rho_{0}[$ in the RJS model. The chosen value $N=1637$ is identical to the number of considered companies but the RJS curve for $N=10000$ is identical on graphical precision (with a slightly modified value for $\varepsilon$).
@@
Figure S10: Comparison of the Lorenz curve for the 1637 companies of the London stock exchange FTSE at 31 December 2024 (black; data from Ref. [34]) with the corresponding curve for the RJS model (red curve; $N=1637$) at same Gini coefficient $G=0.9126$ obtained for $\varepsilon=0.04387$. The top (bottom) panel shows the full range $h\in[0,1]$ (zoomed range $h\in[0.8,1]$).
@@
Fig. S11, compares the Lorenz curve for the Hong Kong stock exchange (2025; data from Ref. [35]) with the RJS model. Here, the Gini coefficient $G=0.9471$ is even higher than for the London stock exchange and the corresponding value $\varepsilon\approx 0.027$ for the RJS model is even smaller. Due to the very high value of $G$ the first probability $\rho_{0}=0.7768$ is even higher (than $\rho_{0}$ for the London stock exchange) indicating a strong RJ condensation with exactly $w=0$ for the larger interval $h\in[0,\rho_{0}[$ in the RJS model. The chosen value $N=2683$ is identical to the number of considered companies but the RJS curve for $N=10000$ is identical on graphical precision (with a slightly modified value for $\varepsilon$).
@@
Figure S11: Comparison of the Lorenz curve for the 2683 companies of the Hong Kong stock exchange at 19 June 2025 (black; data from Ref. [35]) with the corresponding curve for the RJS model (red curve; $N=2683$) at same Gini coefficient $G=0.9471$ obtained for $\varepsilon=0.02651$. The top (bottom) panel shows the full range $h\in[0,1]$ (zoomed range $h\in[0.8,1]$).
@@
Fig. S12, compares the Lorenz curve for the 30 Dow Jones companies (2025; data from Ref. [33]) with the RJS model. Here, the Gini coefficient $G=0.3096$ is very low and the corresponding value $\varepsilon\approx 0.55$ for the RJS model is very high being in the region for $T<0$ with large $|T|$. The value of $G$ is even smaller then $G=1/3$ for the curve $w=h^{2}$ corresponding to the RJS model with $\varepsilon=0.5$ and $|T|\to\infty$. The chosen value $N=30$ is identical to the number of considered companies but despite the modest value of $N$ the RJS curve for $N=10000$ is identical on graphical precision (with a slightly modified value for $\varepsilon$). We mention, that a comparison with the EQI model for a modest value of $E_{0}$ to fit approximately the finite initial slope in the data provides the energy value $\varepsilon\approx 0.48<0.5$ corresponding to the regime of $T>0$ but still with large $|T|$. We note that this case is very special since these 30 companies are certainly not isolated and they constitute a subset of the 504 companies of S&P500 (which are not perfectly isolated as well).
@@
Figure S12: Comparison of the Lorenz curve for the 30 Dow Jones companies of NYSE 2025 (black; data from the site of Ref. [33] taken at June 18, 2025) with the corresponding curve for the RJS model (red curve; $N=30$) at same Gini coefficient $G=0.3096$ obtained for $\varepsilon=0.5528$. The dashed green (black) line represents the curve for $w=h^{2}$ ($w=h$) for the RJS model at $\varepsilon=0.5,T\to\infty$ (perfect equipartition).
For the data of NYSE S&P500 (and Dow Jones companies) the assumption that the system is completely isolated and the total wealth is preserved is somewhat less justified (or not all justified for the Dow Jones case) in comparison to populations of countries or the whole world. We suppose that this fact is the reason for the reduced fraction of poor companies since external fluctuations may increase energies. However, in global the Lorenz curves shown in the section for different stock exchange selections are rather well described by the RJS model but indeed with strongly varying values of $\varepsilon$ according to the different values of the Gini coefficient with either $G\approx 1-2\varepsilon$ (if $\varepsilon\ll 1$) or $G\approx 1/3$ (if $w\approx h^{2}$ as in the Dow Jones case).
V V. Extended RJS model
The comparison of the different data with the RJS model shows that typically the curves of the RJS model have a slower (final) growth rate. Since the latter is proportional to the energy $E_{m}$ one could try an extended model where the energy values grow stronger with $m$. One step in this direction is the DL model which allowed for a considerable improvement as can be seen in the right panel of Fig. 4 and also its zoomed version Fig. S5. Another possibility is to choose an exponential growth of $E_{m}$ but in such a way that still $E_{m}\sim m$ for small $m$. This can be achieved by the formula
$E_{m}=\frac{e^{a\,(m/N)}-1}{a}$ (S.16)
which we call RJE model (RJ extended or RJ exponential model). Here $a$ is an additional parameter of the model in addition to the value of $\varepsilon$ which is now $\varepsilon=E/B$ with bandwidth $B=E_{N-1}\approx{(e^{a}-1)/a}$. In the limit $a\to 0$, we recover simply the RJS model and with increasing values of $a$ the exponential growth of $E_{m}$ becomes more dominant.
The energy spectrum (S.16) corresponds to a density of states:
$\nu(E_{m})=\frac{dm}{dE_{m}}=\frac{d}{dE_{m}}\left(\frac{N\ln(1+aE_{m})}{a}\right)=\frac{N}{1+aE_{m}}$ (S.17)
To determine optimal values for the parameter $a$, we compute a reconstructed spectrum from a given Lorenz curve of some given data set (see next section for a description and more detailed discussion of this spectral reconstruction) and fit the reconstructed spectrum to the function $E_{m}\approx C(e^{a\,(m/N)}-1)$ with two parameters $C$ and $a$. The 2nd parameter $C$ has no importance since one could apply an arbitrary fixed factor on (S.16) without changing the resulting Lorenz curve of the RJE model. This is because the construction procedure of Lorenz curve involves only the ratio $E_{m}/E$ (with $E$ being the average energy) so that the global energy scale (or bandwidth $B$) drops out.
Figure S13: Comparison of the Lorenz curve for the data of US 2019 (black) with the corresponding curves for the RJS model (red curve; $N=10000$) and the RJE model with $a=4.42$ (blue curve; $N=10000$). The rescaled energy values $\varepsilon=0.01233$ (RJE) and $\varepsilon=0.07420$ (RJS) are obtained by matching the Gini coefficient $G=0.8515$. The value of $a$ is obtained by a fit from the reconstructed spectrum. The top (bottom) panel shows the full range $h\in[0,1]$ (zoomed range $h\in[0.8,1]$).
The results for US 2019, UK 2012-2014, World 2010, FTSE 2024 (London stock exchange) and the Hong Kong 2025 stock exchange are shown in Figs. S13S17, in each case with two panels, top for the full range of $h\in[0,1]$ and bottom for the zoomed range $h\in[0.8,1]$. Here, we choose for simplicity the value of $N=10000$ for the curves of both RJE and RJS models (the RJS curves are also shown for comparison). Other values such as $N=1000$ or the given size of the data set, give the same Lorenz curves at graphical precision.
Figure S14: Comparison of the Lorenz curve for the data of UK 2012-2014 (black) with the corresponding curves for the RJS model (red curve; $N=10000$) and the RJE model with $a=2.18$ (blue curve; $N=10000$). The rescaled energy values $\varepsilon=0.1332$ (RJE) and $\varepsilon=0.1996$ (RJS) are obtained by matching the Gini coefficient $G=0.6255$. The value of $a$ is obtained by a fit from the reconstructed spectrum. The top (bottom) panel shows the full range $h\in[0,1]$ (zoomed range $h\in[0.8,1]$).
Figure S15: Comparison of the Lorenz curve for the data of World 2021 (black) with the corresponding curves for the RJS model (red curve; $N=10000$) and the RJE model with $a=5.34$ (blue curve; $N=10000$). The rescaled energy values $\varepsilon=0.008553$ (RJE) and $\varepsilon=0.07911$ (RJS) are obtained by matching the Gini coefficient $G=0.8420$. The value of $a$ is obtained by a fit from the reconstructed spectrum. The top (bottom) panel shows the full range $h\in[0,1]$ (zoomed range $h\in[0.8,1]$).
In all cases, the agreement of the RJE model with the data is significantly better than the RJS model. In particular for HK 2025, the agreement is close to perfect and even in the zoomed panel it is difficult to distinguish the theoretical RJE curve (blue) from the data (black). For the case UK 2012-2014 the original simpler RJS model was already quite good, but also here the RJE model provides a significant improvement. The RJE curves for US 2019 and World 2021 are also very good, nearly as good as the curve for HK 2025. For FTSE 2024 the agreement of the data with the RJE model is a bit less perfect but still clearly better than the RJS model.
Figure S16: Comparison of the Lorenz curve for the data of the London stock exchange FTSE at 31 December 2024 (black; data from Ref. [34]) with the corresponding curves for the RJS model (red curve; $N=10000$) and the RJE model with $a=3.13$ (blue curve; $N=10000$). The rescaled energy values $\varepsilon=0.01346$ (RJE) and $\varepsilon=0.04376$ (RJS) are obtained by matching the Gini coefficient $G=0.9126$. The value of $a$ is obtained by a fit from the reconstructed spectrum. The top (bottom) panel shows the full range $h\in[0,1]$ (zoomed range $h\in[0.8,1]$).
Figure S17: Comparison of the Lorenz curve for the data of the Hong Kong stock exchange at 19 June 2025 (black; data from Ref. [35]) with the corresponding curves for the RJS model (red curve; $N=10000$) and the RJE model with $a=6.83$ (blue curve; $N=10000$). The rescaled energy values $\varepsilon=0.0008381$ (RJE) and $\varepsilon=0.02648$ (RJS) are obtained by matching the Gini coefficient $G=0.9471$. The value of $a$ is obtained by a fit from the reconstructed spectrum. The top (bottom) panel shows the full range $h\in[0,1]$ (zoomed range $h\in[0.8,1]$).
VI VI. Spectral reconstruction procedure
@@
Let us briefly remind the construction of a Lorenz curve, already given in the main part, from a given model spectrum $E_{m}$ with appropriate values of $\mu$, $E$ and $T$ such that the two conditions (S.1) are verified. For this a set of points $(h_{m},w_{m})$ is determined with $h_{0}=w_{0}=0$, $h_{m+1}=h_{m}+\rho_{m}$ and $w_{m+1}=w_{m}+(E_{m}/E)\,\rho_{m}$ for $m=0,\ldots,N-1$ and where
$\rho_{m}=\frac{E-\mu}{N(E_{m}-\mu_{m})}$ (S.18)
@@
(see also (S.2). Then the conditions (S.1) assure that $h_{N}=w_{N}=1$ and the points $(h_{m},w_{m})$ provide for $0\leq m\leq N$ the associated Lorenz curve with both $h_{m},w_{m}\in[0,1]$.
@@
We choose some initial value for $\mu$ (close to zero), start with $m=N-1$ and want first to determine $E_{m}=E_{N-1}$. Here we know the last value $h_{m+1}=1$ at $m=N-1$. Therefore, we can at least approximately compute $E_{m}=w^{\prime}(h_{m+1})$ and then the associated value of $\rho_{m}$ using (S.18). This pair $(E_{m},\rho_{m})$ is yet not very precise since the derivative is taken at the right boundary of the interval $[h_{m},h_{m+1}]$ and we can refine its value by recomputing $E_{m}=w^{\prime}(h_{m+1}-\rho_{m}/2)$ using a small shift with the first approximate value of $\rho_{m}$ (which will then be updated with the more precise value of $E_{m}$ using (S.18)). In principle, one could iterate this refinement step until there is convergence of $(E_{m},\rho_{m})$. However, in our experience the method works best with precisely one refinement step (to ensure later convergence for a good value of $\mu$). Once $\rho_{m}$ is known, we obtain $h_{m}=h_{m+1}-\rho_{m}$. Then we decrease $m$ by 1 and repeat this procedure to compute the next values of $E_{m}$, $\rho_{m}$ and $h_{m}$ at $m=N-2$. This provides a recursion for $m=N-1,N-2,\ldots,1,0$ and three sequences for $E_{m}$, $\rho_{m}$ and $h_{m}$ with decreasing $m$.
@@
Instead of searching numerically the zero of the function $h_{0}(\mu)$, one can also use another more practical method to determine the correct value of $\mu$. For this, one can at the last step $m=0$ manually fix the last density value and compute from (S.18) a new modified value $\tilde{\mu}$ such that the condition $\rho_{0}(\tilde{\mu})=h_{1}$ holds exactly and therefore $h_{0}=0$ is perfectly verified. The modified value $\tilde{\mu}$ can be reinjected in the procedure from the start resulting in a fixed-point iteration for $\mu$ which typically converges quite well and allows also to use the exact initial value $\mu=0$ at the first iteration (which is not a problem since this value is not used in the last step at $m=0$ with $E_{0}=0$). For this method the convergence is typically a bit slower, but more reliable, as compared to the secant method applied to $h_{0}(\mu)$ but the latter fails to converge in cases of bad quality data which influence the computation of $w^{\prime}(h)$. In such a situation the fixed point iteration does not always provide a convergence with high precision as well but still the $\mu$ values stabilize in some small interval (with relative fluctuations $\sim 10^{-3}$ etc.) and any value in this interval can be used to have a nice reconstructed spectrum.
Once the procedure is finished, we can use the obtained spectrum $E_{m}$ to recompute a new appropriate value of $\mu$ and the densities $\rho_{m}$ in the usual way by numerically solving (S.3) with the value $E=1$. In case of good convergence of the procedure this simply confirms the already obtained values of $\mu$ and $\rho_{m}$ but in case of a problematic convergence, this provide a final refinement of $\mu$ and $\rho_{m}$ which will match precisely the spectrum $E_{m}$ with $E=1$ according to (S.1) and (S.2). Using these refined densities, we can finally recompute the Lorenz curve associated to this spectrum in the usual way. This curve matches typically, also in the case of not perfect convergence, very well the original data with numerical errors below $10^{-3}$ (or less).
@@
If the raw data is of good quality, i.e. with support points that lie very accurately on a smooth function, one can also use a combination of rational interpolation (for the region where $w^{\prime}(h)>1$) and polynomial interpolation (for the region with $w^{\prime}(h)<1$) and in both cases with a small number of support points between 3 and 6 which are closest to the value of $h$ for which we want to compute $w^{\prime}(h)$. In both interpolation approaches, one can work out efficient formulas to exactly evaluate the derivative of the interpolation function. However, if the data is of bad quality this procedure may be problematic for $\mu$-convergence and also violate the property that $w^{\prime}(h_{1})<w^{\prime}(h_{2})$ for $h_{1}<h_{2}$ which is crucial to obtain a correctly ordered spectrum (with $E_{m_{1}}\leq E_{m_{2}}$ for $m_{1}<m_{2}$). In such a case, it may be necessary to clean the data by coarse graining them (keeping only 15-20 significant data points) and then recompute a new data set with 500 or 1000 points using high quality interpolation (also rational/polynomial interpolation with 4-5 support points from the reduced set). In particular for the data of UK 2012-2014 from Ref. [31] with a lot of data points but with limited precision this was necessary.

### v3

@@ A.4 Data for companies of stock exchange at New York, London, Hong Kong
First, we present in Fig. A.9 the Lorenz curve for the data of 504 S&P500 companies of the New York Stock Exchange (NYSE) of June 16, 2025 (see Ref. [57]). This Fig. A.9 shows the direct comparison of the Lorenz curve of NYSE and the corresponding RJ thermal distribution of the RJS model (at same Gini value). Here, we use the standard value $N=10000$ for the RJS curve but using a reduced value $N=504$ (as the number of companies) gives the same RJS curve within graphical precision. The quality of agreement with the RJS model is comparable to the cases of US or World in Fig. I.3. We also note the characteristic values: $h=0.191$ at wealth $w=0.02$; $w=0.092$ at $h=0.5$; the wealth of top 10 percent of $h$ companies is 1w(0.9)=0.6021-w(0.9)=0.602 and the wealth of top 1 percent of companies is 1w(0.99)=0.2671-w(0.99)=0.267. Thus we see that there is a small fraction of oligarchic companies that monopolize a big fraction of total wealth. The fraction of poor companies, corresponding to the RJ condensate, is smaller than the fraction of poor households in the US or World cases. We attribute this to the fact that these 504 companies of S&P500 represent only about 80 percent of the total capitalization of US companies.
@@
Figure A.9: Comparison of the Lorenz curve for the S&P500 companies of NYSE 2025 (black; data from Ref. [57]) with the corresponding curve for the RJS model (red curve; $N=10000$) at same Gini coefficient $G=0.692$ obtained for $\varepsilon=0.1582$.
@@
Fig. A.10, compares the Lorenz curve for the London stock exchange (2024; data from Ref. [58]) with the RJS model. Here, the Gini coefficient $G=0.9126$ is higher than for the US and World cases and the corresponding value $\varepsilon\approx 0.044$ for the RJS model is quite small. Due to the high value of $G$ the first probability $\rho_{0}=0.6545$ is very high indicating a strong RJ condensation with exactly $w=0$ for $h\in[0,\rho_{0}[$ in the RJS model. The chosen value $N=1637$ is identical to the number of considered companies but the RJS curve for $N=10000$ is identical on graphical precision (with a slightly modified value for $\varepsilon$).
@@
Figure A.10: Comparison of the Lorenz curve for the 1637 companies of the London stock exchange FTSE at 31 December 2024 (black; data from Ref. [58]) with the corresponding curve for the RJS model (red curve; $N=1637$) at same Gini coefficient $G=0.9126$ obtained for $\varepsilon=0.04387$. The left (right) panel shows the full range $h\in[0,1]$ (zoomed range $h\in[0.8,1]$).
@@
Fig. A.11, compares the Lorenz curve for the Hong Kong stock exchange (2025; data from Ref. [59]) with the RJS model. Here, the Gini coefficient $G=0.9471$ is even higher than for the London stock exchange and the corresponding value $\varepsilon\approx 0.027$ for the RJS model is even smaller. Due to the very high value of $G$ the first probability $\rho_{0}=0.7768$ is even higher (than $\rho_{0}$ for the London stock exchange) indicating a strong RJ condensation with exactly $w=0$ for the larger interval $h\in[0,\rho_{0}[$ in the RJS model. The chosen value $N=2683$ is identical to the number of considered companies but the RJS curve for $N=10000$ is identical on graphical precision (with a slightly modified value for $\varepsilon$).
@@
Figure A.11: Comparison of the Lorenz curve for the 2683 companies of the Hong Kong stock exchange at 19 June 2025 (black; data from Ref. [59]) with the corresponding curve for the RJS model (red curve; $N=2683$) at same Gini coefficient $G=0.9471$ obtained for $\varepsilon=0.02651$. The left (right) panel shows the full range $h\in[0,1]$ (zoomed range $h\in[0.8,1]$).
@@
Fig. A.12, compares the Lorenz curve for the 30 Dow Jones companies (2025; data from Ref. [57]) with the RJS model. Here, the Gini coefficient $G=0.3096$ is very low and the corresponding value $\varepsilon\approx 0.55$ for the RJS model is very high being in the region for $T<0$ with large $|T|$. The value of $G$ is even smaller then $G=1/3$ for the curve $w=h^{2}$ corresponding to the RJS model with $\varepsilon=0.5$ and $|T|\to\infty$. The chosen value $N=30$ is identical to the number of considered companies but despite the modest value of $N$ the RJS curve for $N=10000$ is identical on graphical precision (with a slightly modified value for $\varepsilon$). We mention, that a comparison with the EQI model for a modest value of $E_{0}$ to fit approximately the finite initial slope in the data provides the energy value $\varepsilon\approx 0.48<0.5$ corresponding to the regime of $T>0$ but still with large $|T|$. We note that this case is very special since these 30 companies are certainly not isolated and they constitute a subset of the 504 companies of S&P500 (which are not perfectly isolated as well).
@@
Figure A.12: Comparison of the Lorenz curve for the 30 Dow Jones companies of NYSE 2025 (black; data from the site of Ref. [33] taken at June 18, 2025) with the corresponding curve for the RJS model (red curve; $N=30$) at same Gini coefficient $G=0.3096$ obtained for $\varepsilon=0.5528$. The dashed green (black) line represents the curve for $w=h^{2}$ ($w=h$) for the RJS model at $\varepsilon=0.5,T\to\infty$ (perfect equipartition).
A.5 Spectral reconstruction procedure
@@
Let us briefly remind the construction of a Lorenz curve, already given in the main part, from a given model spectrum $E_{m}$ with appropriate values of $\mu$, $E$ and $T$ such that the two conditions (A.1) are verified. For this a set of points $(h_{m},w_{m})$ is determined with $h_{0}=w_{0}=0$, $h_{m+1}=h_{m}+\rho_{m}$ and $w_{m+1}=w_{m}+(E_{m}/E)\,\rho_{m}$ for $m=0,\ldots,N-1$ and where
$\rho_{m}=\frac{E-\mu}{N(E_{m}-\mu_{m})}$ (A.16)
@@
(see also (A.2)). Then the conditions (A.1) assure that $h_{N}=w_{N}=1$ and the points $(h_{m},w_{m})$ provide for $0\leq m\leq N$ the associated Lorenz curve with both $h_{m},w_{m}\in[0,1]$.
@@
We choose some initial value for $\mu$ (close to zero), start with m=N1m=N-1 and want first to determine $E_{m}=E_{N-1}$. Here we know the last value $h_{m+1}=1$ at m=N1m=N-1. Therefore, we can at least approximately compute $E_{m}=w^{\prime}(h_{m+1})$ and then the associated value of $\rho_{m}$ using (A.16). This pair $(E_{m},\rho_{m})$ is yet not very precise since the derivative is taken at the right boundary of the interval $h_{m},h_{m+1}]$ and we can refine its value by recomputing $E_{m}=w^{\prime}(h_{m+1}-\rho_{m}/2)$ using a small shift with the first approximate value of $\rho_{m}$ (which will then be updated with the more precise value of $E_{m}$ using (A.16)). In principle, one could iterate this refinement step until there is convergence of $(E_{m},\rho_{m})$. However, in our experience the method works best with precisely one refinement step (to ensure later convergence for a good value of $\mu$). Once $\rho_{m}$ is known, we obtain $h_{m}=h_{m+1}-\rho_{m}$. Then we decrease $m$ by 1 and repeat this procedure to compute the next values of $E_{m}$, $\rho_{m}$ and $h_{m}$ at m=N2m=N-2. This provides a recursion for $m=N-1,N-2,\ldots,1,0$ and three sequences for $E_{m}$, $\rho_{m}$ and $h_{m}$ with decreasing $m$.
@@
Instead of searching numerically the zero of the function $h_{0}(\mu)$, one can also use another more practical method to determine the correct value of $\mu$. For this, one can at the last step $m=0$ manually fix the last density value and compute from (A.16) a new modified value $\tilde{\mu}$ such that the condition $\rho_{0}(\tilde{\mu})=h_{1}$ holds exactly and therefore $h_{0}=0$ is perfectly verified. The modified value $\tilde{\mu}$ can be reinjected in the procedure from the start resulting in a fixed-point iteration for $\mu$ which typically converges quite well and allows also to use the exact initial value $\mu=0$ at the first iteration (which is not a problem since this value is not used in the last step at $m=0$ with $E_{0}=0$). For this method the convergence is typically a bit slower, but more reliable, as compared to the secant method applied to $h_{0}(\mu)$ but the latter fails to converge in cases of bad quality data which influence the computation of $w^{\prime}(h)$. In such a situation the fixed point iteration does not always provide a convergence with high precision as well but still the $\mu$ values stabilize in some small interval (with relative fluctuations $\sim 10^{-3}$ etc.) and any value in this interval can be used to have a nice reconstructed spectrum.
Once the procedure is finished, we can use the obtained spectrum $E_{m}$ to recompute a new appropriate value of $\mu$ and the densities $\rho_{m}$ in the usual way by numerically solving (A.3) with the value $E=1$. In case of good convergence of the procedure this simply confirms the already obtained values of $\mu$ and $\rho_{m}$ but in case of a problematic convergence, this provide a final refinement of $\mu$ and $\rho_{m}$ which will match precisely the spectrum $E_{m}$ with $E=1$ according to (A.1) and (A.2). Using these refined densities, we can finally recompute the Lorenz curve associated to this spectrum in the usual way. This curve matches typically, also in the case of not perfect convergence, very well the original data with numerical errors below $10^{-3}$ (or less).
@@
If the raw data is of good quality, i.e. with support points that lie very accurately on a smooth function, one can also use a combination of rational interpolation (for the region where $w^{\prime}(h)>1$) and polynomial interpolation (for the region with $w^{\prime}(h)<1$) and in both cases with a small number of support points between 3 and 6 which are closest to the value of $h$ for which we want to compute $w^{\prime}(h)$. In both interpolation approaches, one can work out efficient formulas to exactly evaluate the derivative of the interpolation function. However, if the data is of bad quality this procedure may be problematic for $\mu$-convergence and also violate the property that $w^{\prime}(h_{1})<w^{\prime}(h_{2})$ for $h_{1}<h_{2}$ which is crucial to obtain a correctly ordered spectrum (with $E_{m_{1}}\leq E_{m_{2}}$ for $m_{1}<m_{2}$). In such a case, it may be necessary to clean the data by coarse-graining them (keeping only 15-20 significant data points) and then recompute a new data set with 500 or 1000 points using high quality interpolation (also rational/polynomial interpolation with 4-5 support points from the reduced set). In particular for the data of UK 2012-2014 from Ref. [31] with a lot of data points but with limited precision this was necessary.
In order, to keep things simple and reliable, we opted for a compromise between rational/polynomial interpolation and a piecewise constant derivative. Without going into too much details, we mention that we computed first discrete derivatives (from good quality data) and applied linear interpolation to obtain a piecewise linear numerical implementation of $w^{\prime}(h)$ which respects that $w^{\prime}(h_{1})<w^{\prime}(h_{2})$ for $h_{1}<h_{2}$ and is still a continuous function. In this approach the support points for $h$ are now in the center of two former supports points (for which the discrete derivative was taken) and also a slight renormalization was applied to assure that the interpolated piecewise linear function satisfies numerically $\int_{0}^{1}w^{\prime}(h)dh=1$, a property which is very important for the reconstruction procedure.
Figure A.13: Rescaled reconstructed spectrum $E_{m}/B$ for the data of Hong Kong 2025 (red data points) versus rescaled level number $x=m/N$. The blue curve shows the curve $E_{m}/B=C\,(e^{ax}-1)$ with $C=0.00160\pm 0.00004$ and $a=6.83\pm 0.03$ obtained from a fit in the interval $x\in[0,0.9]$. The value of $a=6.83$ is used in Fig. I.11 for the blue curve of the RJE model.
Globally, the fit $E_{m}=C(e^{a(m/N)}-1)$ works rather well at least for some reasonable subinterval. For each data set, we perform two fits of this function for the intervals $x\in[0,0.7]$ and $x\in[0,0.9]$ that provides two interesting values of $a$. We inject these two values of the parameter $a$ in the RJE model and determined which $a$ value gives a better agreement for the Lorenz curve (the value of $\varepsilon$ is determined as usual by matching the Gini coefficient to be identical between initial Lorenz curve and the model curve). Figs. I.7I.11 show the resulting RJE curves for 5 of our data sets, already discussed in the previous section, with a very good agreement of the RJE model at the optimal fitted values of $a$.
In Fig. A.13, we show as illustration one example of an reconstructed spectrum for the data set of the stock market Hong Kong 2025 using the value $N=2683$ and the piecewise linear derivative for $w^{\prime}(h)$. The number $N=2683$ represents the number of companies used in the Hong Kong SE data but the precise choice of this value is not very important and the reconstruction procedure works also nicely for $N=1000$ or $N=10000$ for this example. In this case, the shown fit $E_{m}/B=C\,(e^{ax}-1)$ works quite well for the larger interval $x\in[0,0.9]$ and the resulting value of $a=6.83$ provides a nearly perfect Lorenz curve of the RJE model.
We mention that the bandwidth $B$ of the reconstructed spectrum shown in Fig. A.13 is $B=566.6$ and it corresponds to the initial choice $E=1$ to fix the global energy scale such that the rescaled energy of the reconstructed spectrum is $\varepsilon_{rc}=1/B\approx 0.001765$. The value of $\varepsilon_{\rm RJE}$ for the fitted blue curve is slightly modified due to a modified bandwidth of the latter: $\varepsilon_{\rm RJE}\approx\varepsilon_{rc}/[C\,(e^{a}-1)]=0.001193$ which compares to the value $\varepsilon=0.0008381$ given in the caption of Fig. I.11 obtained by matching the Gini coefficient. The slightly different value for $\varepsilon$ is due to the matching of the Gini coefficient and the fact that the fit is far from perfect. Furthermore, also the data points for $x>0.9$ are not very reliable.
Finally, we note that instead of the above reconstruction procedure it is also possible simply to fit the value of parameter $a$ in the spectrum $E_{m}$ of Eq.(I.2) in such a way that the Lorenz curve of the RJE model is closet to the real Lorenz curve which is done by minimizing a suitable metric to measure the distance between two curves (parameter $\varepsilon$ is fixed as usually to match Gini coefficient). We checked that the obtained $a$ values by minimizing different variants of such metrics are rather close to those obtained from the reconstruction procedure. Thus both approaches allows to obtain very good agreement between the model and real Lorenz curves. The advantage of the reconstruction procedure is related to a more physical understanding of the origins of spectrum $E_{m}$.
Appendix B Additional material for Part II
B.1 Netscience network